A person consumes about 2400Cala day.1 Cal = 4186 J, and 1 kWh = 860 Cal. What is the energy in Joules?
Answer #1

2400 * 4186 = 10046400 J PER DAY
Answer #2

2400 Cal = 2400*4186 = 10046400 joules = 2.79 kWh


1 kWh = 860 Cal = 860*4186 = 3599960 joules.